OpenStudy (anonymous):

Find the inverse of f(x) = (x - 8)/3

OpenStudy (anonymous):

first solve for x

OpenStudy (anonymous):

and lets call f(x) y

OpenStudy (anonymous):

y=x-8/3

OpenStudy (anonymous):

ok! now solve for x

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

yes do you sub 8

OpenStudy (anonymous):

noooo not yet. first lets get rid of the denominator. multiply it from bothsides

OpenStudy (anonymous):

f(x)=3x+8

OpenStudy (anonymous):

not really...how about 3y+8=x?

OpenStudy (anonymous):

hello?

OpenStudy (ybarrap):

Another way to try. Let y=f(x). Just replace y with x with and x with y. Then, solve for y and that's your inverse.